OpenStudy (anonymous):

HELP WILL FAN AND GOLD VERY HARD three times a first number minus a second number equals negative forty. The first number plus twice the second number equals negative four a.Define variables and formulate a system of linear equations from this situation. b.What are the numbers

OpenStudy (anonymous):

3n-m=-40 n+2m = -4 does this look right so far?

OpenStudy (anonymous):

hmm, i'm pretty sure

OpenStudy (anonymous):

3 times a number (3n) minus a second number(m) = -40

OpenStudy (anonymous):

can you solve for n on the second equation i gave you and tell me what it equals?

OpenStudy (anonymous):

n=-2?

OpenStudy (anonymous):

n+2m = -4 can you walk me through how you got n=-2?

OpenStudy (anonymous):

HONESTLY I'M SO CONFUSED LOL can you you walk me through how to get n?

OpenStudy (anonymous):

get n alone! n+2m = -4 subtract 2m from both sides!

OpenStudy (anonymous):

so... n =?

OpenStudy (anonymous):

n=-4-2m

OpenStudy (anonymous):

ok! good. now that we know this, lets go back to our first equation 3n-m = -40. can you guess what we do from here or do you need a clue?

OpenStudy (anonymous):

we plug in n?

OpenStudy (anonymous):

wooohoo!

OpenStudy (anonymous):

so 3(-4-2m)=-40m ??

OpenStudy (anonymous):

we have 3(n)- m = -40 put everything that n = inside those parentheses. 3(-4-2m) - m = -40

OpenStudy (anonymous):

and we distribute the 3?

OpenStudy (anonymous):

we do, and then we solve for m. once we have m we plug that in to one of the two equations we came up with and find n and we are finished!